Consider the following half reactions and their standard reduction potential values to answer the following questions.

Cu2+ (aq) + e- Cu+ (aq) E° = 0.15 V
Br2 (l) + 2e- 2Br- (aq) E° = 1.08 V
iii) Calculate the value of E°cell.
iv) Calculate the value ΔG°
v) What is the value of Kc at 25°C?
1
Expert's answer
2020-06-04T10:30:38-0400

iii) E cell = E reduced - E oxidised = 1.08 - 0.15 = 0.93 V

iv) ∆G = - nFE

n = 2e-

∆G = - 2 * 96500 * 0.93 = -179490 J

v) ∆G=-RTlnKc

ln Kc = -∆G / RT

ln Kc = -179490 / 8.314 x 298 = -72.4

Kc = 3.45 x 10-32